OpenStudy (anonymous):

There are 8 people on the debate team. In how many ways can the coach choose 4 members to send to competition?

OpenStudy (anonymous):

you are being asked to compute 8 choose 4 written \[\dbinom{8}{4}\] and computed via \[\dbinom{8}{4}=\frac{8\times 7\times 6\times 5}{4\times 3\times 2}\] cancel first, multiply last or use a calculator that has the appropriate button

OpenStudy (anonymous):

or just use the internet http://www.wolframalpha.com/input/?i=8+choose+4

OpenStudy (anonymous):

doesn't help. :l

OpenStudy (anonymous):

nevermind, it does. ^-^